OpenStudy (anonymous):

8. Solve the equation. Check for extraneous solutions. 6|6 – 4x| = 8x + 4 (1 point) x = 1 x = five over two or x = 2 x = five over two x = five over two or x = 1

OpenStudy (anonymous):

Can I please have help on this

OpenStudy (ranga):

Factor 2 out of the absolute bars: 6 * 2 * |3 - 2x| = 4(2x + 1) divide both sides by 12 |3 - 2x| = 4/12 * (2x + 1) |3 - 2x| = 1/3 * (2x+1) square both sides and solve for x. You will get a quadratic equation with two values for x. Test each x value with the original equation and discard any extraneous solution introduced by squaring. Keep the solution that satisfies the original equation.

OpenStudy (ranga):

Instead of dividing both sides by 12 it may be easier to divide both sides by 4: 3 * |3-2x| = (2x+1) square both sides and solve for x.

OpenStudy (ranga):

A second method to solve the problem WITHOUT squaring: |6 – 4x| is zero when 4x = 6 or x = 1.5 When x < 1.5, 6-4x is positive and therefore, |6-4x| = 6 - 4x When x > 1.5, 6-4x is negative and therefore, |6-4x| = -(6-4x) = 4x - 6 For x < 1.5: 6(6 - 4x) = 8x + 4 36 - 24x = 8x + 4 32x = 32 x = 1. This solution falls under the case of x < 1.5 and therefore x = 1 is a valid solution. For x > 1.5: 6(4x - 6) = 8x + 4 24x - 36 = 8x + 4 16x = 40 x = 40/16 = 5/2 = 2.5. This solution falls under the case of x > 1.5 and therefore x = 2.5 is a valid solution. x = 1 and x = 2.5 are the solutions. You can put both values into the original equation to make sure it satisfies.

OpenStudy (anonymous):

Okay, thank you so much for explaining in detail, this really helped

OpenStudy (ranga):

You are welcome.